Suppose a basketball player is an excellent free throw (shots awarded when a player is fouled) shooter and makes 80% of his free throws (or he has and 80% chance of making a single free throw). Assume that free throw shots are independent of one another. Suppose this player gets to shoot four free throws. Find the probability that he makes four consecutive free throws.
Answer #1

Answer

Given that the probability p = 80/100 = 0.80

Number of trials n = 4

We have to find the probability that he makes four consecutive free throws.

Required probability = .................because each trial is independent, so we can apply multiplication rule

=

= 0.4096 ( 4 decimals)

So, the probability that he makes four consecutive free throws is 0.4096